# Quantum Simulation of Spin-Boson Models with Structured Bath

### v2

@@ Abstract
The spin-boson model, involving spins interacting with a bath of quantum harmonic oscillators, is a widely used representation of open quantum systems. Trapped ions present a natural platform for simulating the quantum dynamics of such models, thanks to the presence of both high quality internal qubit states and the motional modes of the ions that can simulate the relevant quantum degrees of freedom. In our work, we extend the previous body of work that focused on coherent coupling of the spins and bosons to perform quantum simulations with structured dissipative baths using the motional states of trapped ions. We demonstrate the capability for adjusting the bath’s temperature and continuous spectral density by adding randomness to fully programmable control parameters. Subsequently, we simulate the dynamics of various spin-boson models with noise spectral densities constructed from coupling to several dissipative harmonic oscillator modes. The experimental outcomes closely align with theoretical predictions, indicating successful simulation of open quantum systems using a trapped-ion system.
@@
Interaction of a quantum system with its environment in an open dissipative system can significantly influence the dynamics of the system. A paradigmatic model of non-Markovian open quantum systems that captures the quantum nature of both the system and the environment is the spin-boson model Leggett et al. (1987), which is ubiquitous in the study of condensed-matter physics Caldeira and Leggett (1983); Weiss et al. (1987); Schindler et al. (2022), chemical reactions Garg et al. (1985); Gilmore and McKenzie (2005), collective modes in atomic-photonic systems Strack and Sachdev (2011); Bhaseen et al. (2012), and biological light-harvesting complexes Huelga and Plenio (2013); Jang and Mennucci (2018). Experimentally realizing the dynamics of spin-boson models in a programmable fashion is an interesting challenge for quantum simulators Mostame et al. (2012); Leppäkangas et al. (2018); Magazzù et al. (2018); Kim et al. (2022).
@@
Trapped ions provide a highly controllable system of spins and motional modes, and thus are a natural platform for simulating spin-boson models. The bosonic bath can be simulated using either a large number ($\gtrsim 50$) of motional modes Porras et al. (2008) or a handful of dissipative motional modes Lemmer et al. (2018); Schlawin et al. (2021); MacDonell et al. (2021); Kang et al. (2024). While significant experimental progress has been made in controlling the qubits and motional modes together Jia et al. (2022); Chen et al. (2023a), past works in this direction are limited to simulating a spin coherently coupled to up to two bosonic modes Gorman et al. (2018); MacDonell et al. (2023); Whitlow et al. (2023); Valahu et al. (2023). These previous studies did not consider the effects of dissipation, which is often described by the coupling of a system to a continuum of bath modes, that can determine the system’s long-term behavior such as reaching a steady state.
In this work, we perform quantum simulations of spin-boson models with continuous spectral densities using trapped ions and their motional modes. Specifically, we simulate spectral densities composed of several Lorentzian peaks, which is characteristic of baths in light-harvesting protein complexes Olbrich et al. (2011); Chandrasekaran et al. (2015); Lee and Coker (2016); Kim et al. (2018); Tamascelli et al. (2019). Furthermore, we simulate the sub-Ohmic, Ohmic, and super-Ohmic spin-boson models Leggett et al. (1987), which are ubiquitous in various open quantum systems ranging from nanomechanical devices Seoanez et al. (2007); Galland et al. (2008); Chen et al. (2011) and superconducting circuits Magazzù et al. (2018); Thorwart et al. (2004); Kaur et al. (2021); Gulácsi and Burkard (2023) to molecules in which proton-transfer reactions occur Cukier and Morillo (1989); Ohta et al. (2006); Craig et al. (2007). Finally, we simulate the vibrationally-assisted energy transfer (VAET) Gorman et al. (2018) in a simple molecular model where dephasing is added to the vibrational mode. The various spin-state population dynamics, which display behaviors such as damped oscillations, coherent revivals, and relaxation to a steady state, are accurately captured by our quantum-simulation results. We note that a recent study Wang et al. (2024) simulated a spin dissipatively coupled to many modes, albeit with limited tunability.
Capturing the dissipative behavior of the spin-boson model, as well as tuning the bath’s temperature, is enabled by adding randomness to the control parameters, such as the laser frequency and phase. Remarkably, the impact of these control signals is directly analogous to motional dephasing, the leading source of noise in our experimental setup Kang et al. (2023a). This serves as a lower limit for the level of noise that can be introduced in our quantum simulation process. While the quantum simulations presented here are limited to the weak spin-bath coupling regime, our demonstration of adding or using noise Guimarães et al. (2023); Leppäkangas et al. (2023) in analog quantum simulations is an important ingredient for a potential quantum advantage in simulating open quantum system models that are intractable using classical simulation methods MacDonell et al. (2021); Kang et al. (2024).
Results
Models
@@ Results
A spin-boson model with a structured bath is often characterized by a spectral, density $J(\omega)$, composed of multiple Lorentzian peaks, each of which can be represented by a noisy harmonic oscillator Lemmer et al. (2018). Specifically, we consider a spin coupled to dephased oscillators (hereafter referred to as the dephased spin-oscillator model). The single-oscillator case is described by the Hamiltonian
$\hat{H}_{SO}=\hat{H}_{S}+\frac{\kappa}{2}\hat{\sigma}_{Z}\otimes(\hat{b}+\hat{b}^{\dagger})+\nu\hat{b}^{\dagger}\hat{b}$ (2)
@@
and a single Lindblad operator $\hat{L}_{1}=\sqrt{\Gamma}\hat{b}^{\dagger}\hat{b}$, where $\hat{b}$ is the annihilation operator of the oscillator, such that the composite state $\rho$ of spin and oscillators follows the Lindblad master equation $\dot{\rho}=-i[\hat{H}_{SO},\rho]+\sum_{j}(\hat{L}_{j}\rho\hat{L}_{j}^{\dagger}-\frac{1}{2}\{\hat{L}_{j}^{\dagger}\hat{L}_{j},\rho\})$ Lindblad (1976). In a reasonable regime of parameters ($\Gamma<\nu/2$, $\beta\ll 2\pi(\Gamma/2)^{-1}$) Lemmer et al. (2018), this bath of dephased oscillators is assigned a spectral density composed of Lorentzian peaks sup
$J_{\rm Lo}(\omega)=\kappa^{2}\Big{(}\frac{\Gamma/2}{(\Gamma/2)^{2}+(\omega-\nu)^{2}}-\frac{\Gamma/2}{(\Gamma/2)^{2}+(\omega+\nu)^{2}}\Big{)},$ (3)
@@
where we assume that the oscillator is initially in the thermal state with average phonon number $\bar{n}(\nu)$.
@@
The spectral density of the bath determines the dynamics of the spin to leading order in $\kappa T$, where $T$ is the evolution time Breuer and Petruccione (2002); sup . Therefore, the spin dynamics of the dephased spin-oscillator model approximately match those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ in the weak-$\kappa$ regime, which extends straightforwardly to the model with multiple oscillators. Of particular interest is the unbiased ($\epsilon=0$) spin-boson model Leggett et al. (1987); Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018), where the equilibrium donor-state population is $1/2$ for both dephased spin-oscillator models and spin-boson models. Thus, for weak spin-bath coupling, the population dynamics of the two models show an excellent match at all times; the match in short (long) time scale is achieved due to matching spectral densities (equilibrium populations). Unless specified otherwise, we use $\epsilon=0$; results for $\epsilon\neq 0$ can be found in Fig. 5.
@@
A more commonly considered model is a spin coupled to damped oscillators (hereafter referred to as the damped spin-oscillator model); for example, the single-oscillator case is described by the Hamiltonian $\hat{H}_{SO}$ and a pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma(\bar{n}(\nu)+1)}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma\bar{n}(\nu)}\hat{b}^{\dagger}$ Lemmer et al. (2018); Schlawin et al. (2021). For this model, a stronger result holds: the spin dynamics are non-perturbatively equivalent to those of the spin-boson model with the same spectral density, up to all orders in $\kappa T$ Tamascelli et al. (2018); sup . Damped oscillators can be realized in trapped ions using sympathetic cooling on a chain of ions with multiple atomic species or isotopes Lemmer et al. (2018); Schlawin et al. (2021). In this work, we focus on dephased oscillators to represent the structured bath, which is readily performed using programmable randomized control operations on a single species of trapped ions.
Single oscillator model
The dephased spin-oscillator model can be simulated using trapped ions by applying spin-dependent kick (SDK) operations, which induce the coupling between the trapped-ion qubit and the motional mode (see Methods for details). By adding randomness to the control parameters of the SDK operations, we can also prepare the thermal state and simulate the dephasing of the motional mode, by averaging the results over many random trials. The key idea is that noisy (heated or dephased) evolution, described by the Lindblad master equation, is equivalent to an average of many coherent evolutions, each subject to a stochastic Hamiltonian Cai and Barthel (2013); Chenu et al. (2017).
@@
Figure 1: Demonstration of the tunability of the bath’s temperature and spectral linewidth. (a) Dynamics of the coherent spin-oscillator model $\hat{H}_{SO}$ with various values of average phonon number $\bar{n}$ of the oscillator’s initial state. (b) Dynamics of the spin-boson model with $\bar{n}=0$, $J(\omega)=J_{\rm Lo}(\omega)$ and various values of $\Gamma$. For both (a) and (b), $\epsilon=0$, $\nu=\Delta$, and $\kappa=0.1\Delta$. Lines (dots) represent theoretical predictions (experimental data) and error bars denote the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times. (c) Revival amplitude versus $\bar{n}$. Predictions of the original (solid line) model and the noise-aware (dashed line) model are compared with the experimental results (circles). (d) Fitted values of $\bar{n}$, where the experimental data are fitted to the predictions of the noise-aware model. Solid line depicts $y=x+\bar{n}_{0}$, where $\bar{n}_{0}=0.036$ is the expected offset in the phonon number due to imperfect cooling. (e) Fitted values of $\Gamma$, where the experimental data are fitted to the theoretical predictions with $\bar{n}=\bar{n}_{0}$. Solid line depicts $y=x$. In (d) and (e), the blue (brown) shaded region represents the expected uncertainty of the respective value when the populations are averaged over 20 (200) random trials. (f) Lorentzian spectral densities for various target (solid) and fitted (dashed) values of $\Gamma$. The shaded region represents the expected uncertainty of $\Gamma$ when averaged over 20 random trials.
@@
We first demonstrate the programmability of the oscillator bath’s temperature, quantified by the average phonon number $\bar{n}$ of the bath oscillator’s initial state. Controlled amounts of resonant SDK operations with stochastically varying phase are applied to the ion in the motional ground state, which prepares the thermal state when averaged over many random trials (see Methods for details). Figure 1a illustrates the simulated coherent time evolution with respect to $\hat{H}_{SO}$, for various values of $\bar{n}$. The dynamics exhibit coherent oscillations with an envelope of slow beat-note that signifies resonant ($\nu=\Delta$) energy exchange between the system and the bath oscillator Eberly et al. (1980). Our initial thermal state is prepared as an ensemble of randomly displaced coherent states. A thermal state of larger $\bar{n}$ is a mixture of more “scattered” coherent states, leading to reduced beat-note amplitude at $t\approx 2\pi/\Delta\times 10.5$. We see a notable agreement between theoretical predictions (solid lines) and experimental results (dots).
@@
We compare our experimental results with the predictions of noise-aware models, reflecting the noise processes in the experimental setup. We use spin-boson models with initial phonon number $\bar{n}+\bar{n}_{0}$ and Lorentzian spectral density with full width at half maximum (FWHM) $\Gamma$ [see Eq. (3)], where the values $\bar{n}_{0}=0.036$ and $\Gamma=0.0022\Delta$ are extracted from separate experiments to characterize imperfect cooling and finite motional coherence time, respectively. The experimentally measured revival amplitude (Fig. 1c) and fitted $\bar{n}$ (Fig. 1d) match well with the predictions of noise-aware models. This shows that well-characterized experimental noise can be used as part of the dissipation model, and thus is not always a bottleneck for the accuracy of our simulations.
Next, we show the tunability of the dephasing strength, which corresponds to the linewidth of the spin-boson model’s spectral density. Dephasing is implemented by adding random frequency offsets to the SDK operations that simulate the time evolution (see Methods for details). This mimics the random fluctuations of the motional-mode frequency, following the stochastic model of motional dephasing noise in trapped-ion systems Wang et al. (2020).
@@
Figure 1b illustrates the dynamics for various values of $\Gamma$ between 0 and $0.4\Delta$, which exhibit a phase transition from underdamped to overdamped dynamics as $\Gamma$ is increased (critical damping condition at $\Gamma\approx 0.18\Delta$). As the damping is increased beyond the critical point, the beat-note that signifies the excitation transfer from the spin to boson back to spin is suppressed, and there is a monotonic decrease in the oscillation of the donor population in the overdamped regime. We notice a $\pi$ phase flip in the oscillation of the donor population in the beat-note in the underdamped case (see, for example, the phase difference between the underdamped and overdamped case at $t\approx 2\pi/\Delta\times 10$). These features match very well with the theoretical predictions of spin-boson models.
@@
The uncertainty of $\bar{n}$ and $\Gamma$ realized in the experiment can be reduced by performing a larger number of random trials, each trial using a different set of random parameters. The shaded areas in Fig. 1d and e show the standard deviation of fitted values for $\bar{n}$ and $\Gamma$, obtained by averaging the results over either 20 (pink) or 200 (blue) random trials. Our experiments are limited to 20 random trials due to the compiling delay required for each set of control parameters, and each trial is repeated 100 times to obtain the expectation value for the donor population. The experimental data is consistent with the standard deviation for 20 trials. We expect to readily increase the number of random trials with fewer repetitions each in future experiments with faster compilation tools Dalvi et al. (2023); sup .
@@
We can use multiple oscillator modes to engineer a target spectral density structure for the oscillator bath. We simulate a bath of spectral density composed of up to 3 Lorentzian peaks, each represented by a motional mode of the ion chain. The spin-boson model with $J(\omega)$ in Fig. 2b is simulated using the SDK operations over three motional modes. Fig. 2a shows the frequency of the laser pulses used to drive the SDK, defined as the detuning of laser beams from the carrier transition frequency. The 1st, 3rd, and 5th motional modes are used, and the detuning of the laser frequency from the motional mode determines the frequency at which the bath spectral density is considered in Figure 2b. Here, we set the electronic coupling strength $\Delta=500$ cm-1 sup ; Chenu and Scholes (2015), such that at temperature 77 K, $\bar{n}\approx 0.1$ for the near-resonance vibrational modes.
Figure 2c shows the results for simulating spin-boson models with spectral densities composed of 2 and 3 Lorentzian peaks, respectively. The theoretical predictions are obtained using the time-dependent density-matrix renormalization group algorithm White and Feiguin (2004); Prior et al. (2010) in the interaction picture Nuomin et al. (2022). The donor population features underdamped coherent oscillations, with contributions from all three modes. The measured population closely follows the theoretical predictions.
The coherence time of the motional modes in our setup of $\sim 5.18(67)$ ms is the leading source of error Wang et al. (2020); Kang et al. (2023a), and is comparable to the experimental time for simulating the time evolution of 2-mode (1.63 ms) and 3-mode (2.56 ms) cases. For these simulations, the contribution of the motional decoherence in the experiment is identical to the random kicks we apply in the spin-oscillator model to simulate the bath, and can be incorporated as the source of dephasing in our simulation with proper calibration Kang et al. (2023a); Milne et al. (2021); Keller et al. (2021).
Figure 2: Simulations of molecular energy transfer in model structured baths. (a) Dashed lines represent the radial motional modes strongly coupled by the Raman beam. Black (grey) lines denote the modes with non-zero (zero) coupling strength to the target ion. The three solid lines (shaded regions) depict the mean value (standard deviation) of the randomized pulse frequencies. (b) Simulated spectral density of the spin-boson model’s bath, where the electronic coupling strength $\Delta=500\>\text{cm}^{-1}$. The dashed curves depict the Lorentzian peak generated by the pulses of corresponding colors in (a). The orange (brown) solid curves illustrate the spectral density summed over two (three) peaks. (c) Time evolution of the spin-boson models with $J(\omega)$ in (b). The bath temperature is 77 K such that $\bar{n}\approx 0.1$, which is set for trapped ions’ motional modes using imperfect sideband cooling. Curves (dots) represent theoretical predictions (experimental data). Error bars denote the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.

### v3

@@ Abstract
The spin-boson model, involving spins interacting with a bath of quantum harmonic oscillators, is a widely used representation of open quantum systems that describe many dissipative processes in physical, chemical and biological systems. Trapped ions present an ideal platform for simulating the quantum dynamics of such models, by accessing both the high-quality internal qubit states and the motional modes of the ions for spins and bosons, respectively. We demonstrate a fully programmable method to simulate dissipative dynamics of spin-boson models using a chain of trapped ions, where the initial temperature and the spectral densities of the boson bath are engineered by controlling the state of the motional modes and their coupling with qubit states. Our method provides a versatile and precise experimental tool for studying open quantum systems.
A new method to simulate open quantum systems made of spins and harmonic oscillators where the bath properties are engineered.
@@
Interaction of a quantum system with its environment in an open dissipative setting can significantly influence its dynamics. A paradigmatic model of non-Markovian open quantum systems that captures the quantum nature of both the system and the environment is the spin-boson model Leggett et al. (1987), which is ubiquitous in the study of condensed-matter physics Caldeira and Leggett (1983); Weiss et al. (1987), chemical reactions Garg et al. (1985); Gilmore and McKenzie (2005), collective modes in atomic-photonic systems Strack and Sachdev (2011); Bhaseen et al. (2012), and biological light-harvesting complexes Huelga and Plenio (2013); Jang and Mennucci (2018). Experimentally realizing the dynamics of spin-boson models in a programmable fashion is an exciting challenge Mostame et al. (2012); Leppäkangas et al. (2018); Magazzù et al. (2018); Kim et al. (2022), which can lead to new quantum simulation approaches that reach beyond classically tractable limit.
@@
Trapped ions provide a highly controllable system of spins and bosonic motional modes, and thus are a natural platform for simulating spin-boson models. The bosonic bath can be simulated using either a large number ($\gtrsim 50$) of motional modes Porras et al. (2008) or a handful of dissipative motional modes Lemmer et al. (2018); Schlawin et al. (2021); MacDonell et al. (2021); Kang et al. (2024). While significant experimental progress has been made in controlling the qubits and motional modes together Chen et al. (2023a); Flühmann and Home (2020); Jia et al. (2022), past works in this direction are limited to simulating a spin coherently coupled to up to two bosonic modes Whitlow et al. (2023); Valahu et al. (2023); Gorman et al. (2018); MacDonell et al. (2023). These studies could not consider the effects of dissipation often described by coupling of the system to a continuum of bath modes, which determines the system’s long-term behavior such as reaching a steady state.
Dissipative couplings come in two flavors: damping is when the quantum system exchanges energy with the bath (inelastic coupling) and eventually reaches thermal equilibrium with the bath, while dephasing is when the phase of the quantum system is randomized by interaction with the bath without energy exchange (elastic coupling). In this work, we perform quantum simulations of spin-boson models with programmable spectral densities using a 7-ion chain and up to 3 of its motional modes. The target spectral densities of the bath are decomposed into several Lorentzian peaks, which is characteristic of baths in real-world systems such as light-harvesting protein complexes Lee and Coker (2016); Kim et al. (2018). Our approach can readily simulate the dephasing model of dissipation, by adding randomness to the control parameters of the laser that drives the spin-boson interaction, such as frequency and phase. Using this technique, we study the dynamics of classic spin-boson models Leggett et al. (1987) and the vibration-assisted energy transfer (VAET) process Gorman et al. (2018). Our demonstration to engineer spectral density of the bath in analog quantum simulations is an important ingredient for achieving quantum advantage in simulating open quantum system models that are intractable using classical methods MacDonell et al. (2021); Kang et al. (2024).
We note that a recent study Wang et al. (2024) simulated a spin dissipatively coupled to many modes, albeit with limited tunability.
Dephased spin-oscillator model
@@
We decompose the spectral density $J(\omega)$ of a structured bath into a sum of multiple Lorentzian peaks, each of which can be represented by a dissipative harmonic oscillator Lemmer et al. (2018). In this work, we consider a spin coupled to several oscillators subject to constant dephasing (“dephased” spin-oscillator model). This “discrete” oscillator model is described by the Hamiltonian
$\hat{H}_{D}=\hat{H}_{S}+\sum_{l}\frac{\kappa_{l}}{2}\hat{\sigma}_{Z}\otimes(\hat{b}_{l}+\hat{b}_{l}^{\dagger})+\sum_{l}\nu_{l}\hat{b}_{l}^{\dagger}\hat{b}_{l}$ (2)
@@
and corresponding Lindblad operators $\hat{L}_{l}=\sqrt{\Gamma}_{l}\hat{b}_{l}^{\dagger}\hat{b}_{l}$, where $\kappa_{l}$ represents the coupling strength between the spin and $l$-th oscillator, $\nu_{l}$ and $\Gamma_{l}$ denote the frequency and the dephasing rate of the $l$-th oscillator, respectively, and $\hat{b}_{l}$ is the annihilation operator of the $l$-th oscillator. The composite state $\rho$ of spin and oscillators follows the Lindblad master equation $\dot{\rho}=-i[\hat{H}_{D},\rho]+\sum_{l}(\hat{L}_{l}\rho\hat{L}_{l}^{\dagger}-\frac{1}{2}\{\hat{L}_{l}^{\dagger}\hat{L}_{l},\rho\})$. In a reasonable regime of parameters ($\Gamma_{l}<\nu_{l}/2$, $\beta\ll 2\pi(\Gamma_{l}/2)^{-1}$) Lemmer et al. (2018), this bath of dephased oscillators is assigned a spectral density composed of Lorentzian peaks sup
$J_{\rm Lo}(\omega)=\sum_{l}\kappa_{l}^{2}\Big{(}\frac{\Gamma_{l}/2}{(\Gamma_{l}/2)^{2}+(\omega-\nu_{l})^{2}}-\frac{\Gamma_{l}/2}{(\Gamma_{l}/2)^{2}+(\omega+\nu_{l})^{2}}\Big{)},$ (3)
@@
where we assume that each oscillator is initially in the thermal state with average phonon number $\bar{n}(\nu_{l})$.
@@
The spectral density of the bath determines the dynamics of the spin to leading order in $\bar{\kappa}T$, where $\bar{\kappa}$ is the average coupling strength and $T$ is the evolution time sup . Therefore, the spin dynamics of the dephased spin-oscillator model approximately match those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ in the weak-$\bar{\kappa}$ regime. In the unbiased ($\epsilon=0$) spin-boson model Leggett et al. (1987); Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018), the equilibrium donor-state population is $1/2$ for both models, and we expect the population dynamics of the two models show an excellent match at all times for weak spin-bath coupling. For the case of biased spin-boson models in the stronger coupling regime ($\epsilon\neq 0$, such as discussed in the VAET case), the dynamics of the dephased spin-oscillator model we simulate can deviate significantly from the traditional spin-boson model.
Figure 1: Schematic of the experimental setup. The ions are confined in a micro-fabricated surface trap. The spin states are encoded as the hyperfine clock states (qubit states) and the bosonic bath modes are encoded as the collective radial motional modes of the ion chain. We strategically use the center ion and the three symmetric modes (except for the center-of-mass mode, as shown in the upper left corner) in the experiment to minimize cross-mode couplings sup . The laser frequencies for the SDK operations are determined by the frequency difference between the two Raman beams, and are depicted by the red and blue lines on the frequency axis detuned by $\delta_{m}$ from the mode frequency $\omega_{m}$, with the color gradients indicating the random frequencies used to implement the decoherence simulation.
@@
A more common model for dissipation is a spin coupled to oscillators subject to constant damping (“damped” spin-oscillator model) described by a pair of Lindblad operators $\hat{L}_{l1}=\sqrt{\Gamma_{l}(\bar{n}(\nu_{l})+1)}\hat{b}_{l}$ and $\hat{L}_{l2}=\sqrt{\Gamma_{l}\bar{n}(\nu_{l})}\hat{b}_{l}^{\dagger}$ Lemmer et al. (2018); Schlawin et al. (2021), where $\Gamma_{l}$ here is the damping rate of the $l$-th oscillator. A stronger result holds for this model: the spin dynamics are non-perturbatively equivalent to those of the spin-boson model with the same spectral density, up to all orders in $\bar{\kappa}T$ Tamascelli et al. (2018); sup . Damped oscillators can be realized in trapped ions using sympathetic cooling on a chain of ions with multiple atomic species or isotopes Lemmer et al. (2018); Schlawin et al. (2021). The dephased oscillator model developed in this work expands the ability to simulate a broader set of bath models not experimentally accessible before.
Experimental Implementation
@@
Figure 2: Demonstration of the tunability of the bath’s initial temperature and spectral linewidth. (a) Dynamics of the coherent spin-single oscillator model $\hat{H}_{D}$ with various values of average phonon number $\bar{n}$ for the oscillator’s initial state. (b) Dynamics of the model with $\bar{n}=0$, $J(\omega)=J_{\rm Lo}(\omega)$ (single peak) and various values of $\Gamma_{1}\equiv\Gamma$. For both (a) and (b), $\epsilon=0$, $\nu_{1}=\Delta$, and $\kappa_{1}=0.1\Delta$. Lines and dots represent theoretical predictions and experimental data, respectively. The error bars (size comparable to the symbols in these two plots) denote measured standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times. (c) Revival amplitude versus $\bar{n}$. Predictions of the original (solid line) model and the noise-aware (dashed line) model are compared with the experimental results (circles). (d) Measured values of $\bar{n}$ fitted to the predictions of the noise-aware model. Solid line depicts $y=x+\bar{n}_{0}$, where $\bar{n}_{0}=0.036$ is the expected offset in the phonon number due to imperfect cooling. (e) Measured values of $\Gamma$ fitted to the theoretical predictions with $\bar{n}=\bar{n}_{0}$. Solid line depicts $y=x$. In (d) and (e), the blue and brown shaded regions represent the expected uncertainty of the value when the populations are averaged over 20 and 200 random trials, respectively. (f) Lorentzian spectral densities for various target (solid) and fitted (dashed) values of $\Gamma$. The shaded region represents the expected uncertainty of $\Gamma$ when averaged over 20 random trials.
The simulations of spin-boson models are performed using a linear chain of seven 171Yb+ ions (Figure 1). Two hyperfine internal states (|$\ket{0}$ and |$\ket{1}$) are used to represent the spin (or qubit). The ions sit 68 $\mu$m above the surface of the trap Revelle (2020), and the heating rate and decoherence time of the zig-zag motional mode are $3.6(3)$ quanta/s and $5.2(7)$ ms, respectively. The other motional modes used have similar magnitudes. Details of our experimental setup can be found in Ref. Wang et al. (2020). Using Doppler, electromagnetically-induced transparency (EIT), and sideband cooling techniques, the zig-zag motional mode can be cooled to near ground state with $\bar{n}=0.036(16)$. Standard qubit manipulation techniques can be used to initialize and measure the qubits, and single-qubit operations are driven by stimulated Raman transitions. The spin-oscillator coupling is simulated using the spin-dependent kick (SDK) operation induced by simultaneous application of blue- and red-sideband transitions sup .
The evolution of $\hat{H}_{D}$ in Eq. (2) is mapped to a sequence of single-qubit rotations and SDK operations via Trotterization. We apply time evolution operator over a short time interval corresponding to each term in the Hamiltonian in interaction picture, and repeat them over many time intervals to simulate the evolution dynamics sup . This approach can be readily extended to simulating more complicated molecular models consisting of many electronic states and bath modes Sun et al. (2023); Kang et al. (2024).
The dephased spin-oscillator model can be simulated using trapped ions by applying the SDK operations, which induce the coupling between the qubit and the motional mode. By adding randomness to the control parameters of the SDK operations and averaging the results over many random trials, we can implement dissipative processes such as preparing the thermal state and simulating the dephasing of the motional mode. The key idea is that certain dissipative evolutions described by the Lindblad master equation are equivalent to an average of many coherent evolutions, each subject to a stochastic Hamiltonian Chenu et al. (2017). Details of the procedure is described in Supplementary Material sup .
Single-oscillator case
@@
We first consider a single-oscillator case ($l=1$), and demonstrate the programmability of the bath’s initial temperature quantified by the oscillator’s average phonon number $\bar{n}$. Controlled amounts of resonant SDK operations with stochastically varying phase are applied to the ion prepared near the motional ground state, which prepares the thermal state equivalent to an ensemble of randomly displaced coherent states sup . Figure 2a illustrates the simulated time evolution of the donor state population when it is coupled to such initial bath state, for various values of $\bar{n}$. The dynamics exhibit coherent oscillations with an envelope of slow collapse and revival (“beat-note”) that signifies resonant ($\nu_{1}=\Delta$) energy exchange between the system and the bath oscillator Eberly et al. (1980). A thermal state of larger $\bar{n}$ is a mixture of more “scattered” coherent states, leading to reduced beat-note amplitude at $t\approx 2\pi/\Delta\times 10.5$.
@@
We compare our experimental results with the predictions of noise-aware models, reflecting the noise processes in the experimental setup. We fit the experimental data using an oscillator model with initial phonon number $\bar{n}+\bar{n}_{0}$ and Lorentzian spectral density with full width at half maximum (FWHM) $\Gamma_{1}\equiv\Gamma$ [Eq. (3)], where the values $\bar{n}_{0}=0.036$ and $\Gamma=0.0022\Delta$ are extracted from independent measurements characterizing imperfect cooling and finite motional coherence time, respectively. The experimentally measured revival amplitude (Fig. 2c) and fitted $\bar{n}$ (Fig. 2d) match well with the predictions of noise-aware models. This shows that well-characterized experimental noise can serve as the baseline values for the thermal excitation and dephasing rate in the open quantum system model, and thus is not always a bottleneck for the accuracy of our simulations.
@@
Next, we show the tunability of the dephasing rate $\Gamma$, which corresponds to the linewidth of the spin-boson model’s spectral density. Dephasing is implemented by adding random frequency offsets to the SDK operations that simulate the time evolution sup . This mimics the random fluctuations of the motional-mode frequency, following the stochastic model of motional dephasing noise in trapped-ion systems Wang et al. (2020). Figure 2b illustrates the dynamics for various values of $\Gamma$ between 0 and $0.4\Delta$, which exhibit a crossover from underdamped to overdamped dynamics as $\Gamma$ is increased (critical damping condition at $\Gamma\approx 0.18\Delta$). As the damping is increased beyond the critical point, the beat-note that signifies the excitation transfer from the spin to boson back to spin is suppressed, and there is a monotonic decrease in the oscillation of the donor population in the overdamped regime. We notice a $\pi$ phase flip in the oscillation of the donor population in the beat-note in the underdamped case, signified by the phase difference between the underdamped and overdamped cases at $t\approx 2\pi/\Delta\times 10$.
Figure 3: Simulations of molecular energy transfer in model structured baths. (a) Simulated spectral density of the spin-boson model’s bath, where the electronic coupling strength $\Delta=500\>\text{cm}^{-1}$. The orange and brown solid curves illustrate the spectral densities summed over two and three peaks, respectively. (b) Dashed lines represent the radial motional modes that are coupled by the Raman beam. Black or grey dashed lines represent modes with non-zero or zero coupling strength to the target ion, respectively. The three solid lines and the shaded regions depict the mean value and the standard deviation of the randomized pulse frequencies, and the color corresponds to the Lorentzian peak generating the final spectral density in (a). (c) Time evolution of the spin-boson models with $J(\omega)$ in (a). The bath temperature is 77 K such that $\bar{n}\approx 0.1$, which is set for trapped ions’ motional modes using imperfect sideband cooling. Curves and dots represent theoretical predictions and experimental data, respectively. Error bars are derived as in Fig. 2.
@@
The uncertainty of $\bar{n}$ and $\Gamma$ realized in the experiment can be reduced by performing a larger number of random trials, each trial using a different set of random parameters. The shaded areas in Fig. 2d and e show the standard deviation of fitted values for $\bar{n}$ and $\Gamma$, obtained by numerical simulation of results averaged over 20 (pink) or 200 (blue) random trials. Our experiments are limited to 20 random trials due to the compiling delay required for each set of control parameters, and each trial is repeated 100 times to obtain the expectation value for the donor population. The experimental data is consistent with the standard deviation derived from 20 numerical trials. We expect to readily increase the number of random trials with fewer repetitions each in future experiments with faster compilation tools Dalvi et al. (2023); sup .